Dan is three times as old as Sam. In 5 years Dan will be two times as old as Sam will be in 5 years. How old is each now?
Found 2 solutions by texttutoring, checkley77:
Answer by texttutoring(324)

Let d = Dan's age
Let s = Sam's age
Two Equations:
d=3s
d+5 = 2(s+5)

Now substitute d=3s into the second equation:

3s+5 = 2(s+5)
3s+5 = 2s + 10
s = 5

Now find d:

d=3s
d=3(5)
d=15

Sam is 5 years old, and Dan is 15 years old.

Answer by checkley77(12844)

D=3X NOW.
S=X NOW.
IN 5 YEARS:
(3X+5)=2(X+5)
3X+5=2X+10
3X-2X=10-5
X=5 SAM'S AGE NOW.
D=3*5=15 YEARS OLD NOW.
PROOF:
(3*5+5)=2(5+5)
15+5=2*10
20=20
IN 5 YEARS DAN WILL BE 15+5=20 YEARS OLD..
IN 5 YEARS SAM WILL BE 5+5=10 YEARS OLD.